# GHOST: Solving the Traveling Salesman Problem on Graphs of Convex Setsthanks: Project Page: https://sites.google.com/view/ghost-gcs-tsp

### v1

@@ 27       $\text{OPEN}\leftarrow\text{OPEN}\cup\{n_{c}\}$
Path Unfolding
We present our (abstract-)path-unfolding algorithm, which uses a multi-label A-like (grenouilleau2019multi) best-first search to enumerate all paths unfolded from a given abstract path $\hat{\pi}=(v_{1},v_{2},\ldots,v_{k})$ on $\hat{G}$ in non-decreasing order of their lower-bound costs. Each search state consists of a triplet $p$ and a label $\ell$, indicating progress toward the next endpoint $v_{\ell+1}$. To allow for cyclic and next-best paths to a state, each search node $n$ consists of its path $n.\pi$ (instead of a triplet) and label $n.\ell$. The node also stores the current cost $n.g$ and an estimated final cost $n.f=n.g+h_{\hat{\pi}}(p,n.\ell)$, where $h_{\hat{\pi}}(p,n.\ell)$ is an admissible heuristic underestimating the cost to sequentially visit all remaining endpoints $v_{\ell+1},\ldots,v_{k}$. Fig. 2-(bc) shows that the optimal trajectory conditioned on some $\pi$ may have a smaller cost than that for another $\pi^{\prime}$, even when $\mathcal{L}(\pi)>\mathcal{L}(\pi^{\prime})$. This motivates exhaustively exploring all unfolded paths for $\hat{\pi}$ to guarantee finding the one that yields the optimal trajectory.
Pseudocode (Alg. 1) The search maintains an OPEN list of nodes, prioritized by $n.f$ (Line 1). Nodes are initialized for every triplet $p$ that matches the prefix of $\hat{\pi}$, with their labels set according to how many initial vertices of $\hat{\pi}$ are matched (Lines 1-1). At each step, the node $n$ with lowest $f$ is popped (Line 1). For an open $\hat{\pi}$, if $n.\pi$ ends at $v_{k}$ and $n.\ell=k$, the search outputs $n.\pi$ as the next-best unfolded path (Line 1). For a closed $\hat{\pi}$, if $n.\pi$ start and ends at the same triplet $(\cdot,v_{k},\cdot)$ and $n.\ell=k$, the search outputs $n.\pi$ as the next-best unfolded path (Line 1), with $n.g$ correctly captures the wraparound triplet cost. The search then branches on all triplets $p$ consistent with the suffix of $n.\pi$. Each child node $n_{c}$ increments the label if appending $p$ reaches endpoint $v_{n.\ell+1}$ (Line 1), updates the path cost $n_{c}.g$ by adding $b_{p}$ and sets $n_{c}.f$ as the sum of $g$ and an admissible heuristic value $h_{\hat{\pi}}$ (Line 1). Optionally, nodes with $n_{c}.f>\bar{c}$ are pruned (Line 1), supporting efficient hierarchical search in GHOST. We employ a multi-label heuristic inspired by li2021lifelong; zhong2022optimal:
Lemma 2.
Proof 2.
Consider any open path $\pi=\pi_{v_{1},v_{2}}\circ\ldots\circ\pi_{v_{k-1},v_{k}}$ unfolded from $\hat{\pi}$, where each subpath $\pi_{v_{i},v_{i+1}}=(v_{i}=w^{(i)}_{1},w^{(i)}_{2},\ldots,w^{(i)}_{\mathcal{K}_{i}}=v_{i+1})$, with current state $p=(u,v,w)$ at label $n.\ell$, followed by $p^{\prime}=(v,w,u^{\prime})$. Its cost-to-go can be decomposed as $b_{(v,w,u)}+\mathcal{L}(\pi_{w,v_{\ell+1}})+\sum_{i=\ell+1}^{k-1}\left(b_{(w^{(i)}_{\mathcal{K}_{i}-1},v_{i},w^{(i+1)}_{2})}+\mathcal{L}(\pi_{v_{i},v_{i+1}})\right)$, which is at least $\mathcal{L}(\pi^{*}_{w,v_{\ell+1}})+\sum_{i=\ell+1}^{k-1}\mathcal{L}(\pi^{*}_{v_{i},v_{i+1}})=h_{\hat{\pi}}(p,n.\ell)$. For closed paths, the inequality still holds with an additional non-negative term $b_{(w^{(k-1)}_{\mathcal{K}_{k-1}-1},v_{k},w^{(1)}_{2})}$ on the left. In summary, $h_{\hat{\pi}}$ aggregates the minimum possible subpath costs but does not account for the triplet costs for $p^{\prime}$ and trajectory endpoints in the actual cost-to-go.
@@
While the ILP formulation for TSP (miller1960integer) can be adapted to RTSP with a static pairwise distance metric over vertices, this is insufficient for GCS, where edge costs do not provide meaningful estimates or capture the “volume” of tour endpoint sets. Instead, we leverage the triplet-based distance metric established above and reformulate RTSP over all abstract passage triplets $\hat{p}\in\hat{\mathcal{P}}$. Each $\hat{p}=(u,v,w)\in\hat{\mathcal{P}}$ is formed from every pair of abstract edges $(u,v),(v,w)\in\hat{E}$, with lower-bound cost: $b_{\hat{p}}=\frac{1}{2}\mathcal{L}(\pi^{*}_{u,v})+L(\hat{p})+\frac{1}{2}\mathcal{L}(\pi^{*}_{v,w})$, where
$L(\hat{p})=\begin{cases}b_{(u,v,w)},&\text{if }(u,v),(v,w)\in E\\ \min\limits_{p=(u,v,\cdot)\in\mathcal{P}}b_{p},&\text{if }(u,v)\in E,(v,w)\notin E\\ \min\limits_{p=(v,w,\cdot)\in\mathcal{P}}b_{p},&\text{if }(u,v)\notin E,(v,w)\in E\\ \min\limits_{p=(\cdot,v,\cdot)\in\mathcal{P}}b_{p},&\text{if }(u,v),(v,w)\notin E\end{cases}$
The path cost $\mathcal{L}(\pi^{*}_{u,v})$ for $(u,v)\in E$ and the triplet cost $b_{(u,v,w)}$ are typically zero, unless complex constraints apply on the GCS vertices or edges. We define $\mathcal{L}(\hat{\pi})=\sum_{\hat{p}\text{ along }\hat{\pi}}b_{\hat{p}}$ and show the following theorem:
@@
The second inequality follows Theorem 1. Consider $\hat{\pi}=(v_{1},\ldots,v_{k})$ and an unfolded $\pi=\pi_{v_{1},v_{2}}\circ\ldots\circ\pi_{v_{k-1},v_{k}}$, where each subpath $\pi_{v_{i},v_{i+1}}=(v_{i}=w^{(i)}_{1},w^{(i)}_{2},\ldots,w^{(i)}_{\mathcal{K}_{i}}=v_{i+1})$. We decompose $\mathcal{L}(\pi)=\sum_{i=1}^{k}\mathcal{L}(\pi_{v_{i},v_{i+1}})+\sum_{i=1}^{k}b_{(w^{(i-1)}_{\mathcal{K}_{i-1}-1},v_{i},w^{(i)}_{2})}$, where $v_{k+1}=v_{1}$ (wraparound) and $w^{(0)}_{\mathcal{K}_{0}}=w^{(k-1)}_{\mathcal{K}_{k-1}-1}$ for notation convenience. As each $\mathcal{L}(\pi_{v_{i},v_{i+1}})\geq\mathcal{L}(\pi^{*}_{v_{i},v_{i+1}})$ by definition and $b_{(w^{(i-1)}_{\mathcal{K}_{i-1}-1},v_{i},w^{(i)}_{2})}\geq L_{(v_{i-1},v_{i},v_{i+1})}$ by construction, summing over all i=1,,ki=1,...,k concludes the proof.
@@
$\sum_{\hat{p}\in\hat{\mathcal{P}}^{\text{ind}}_{e}}y_{\hat{p}}=2,$ $\forall(u,v)\in\hat{E}^{+}$ (8)
@@ 25    $\tau\leftarrow$ optimal trajectory conditioned on $\pi$
Properties of GHOST
As in van1999solving; ren2023cbss, the Lawler-Murty procedure at the high level of GHOST guarantees to explore TSP tours in non-decreasing order of their lower-bound costs. Theorem 4 establishes the optimality of GHOST, even when pruning is applied.
Corollary 4.
GHOST returns an optimal GCS-TSP solution.
Proof 4.
This follows from Theorem 5 with $\epsilon=0.0$.
@@
We introduce a bounded-suboptimal variant, $\epsilon$-GHOST, which trades optimality for computational efficiency. Given a parameter $\epsilon\in[0,1)$, we modify GHOST in two ways: On Line 2, the termination condition is relaxed from $c(n^{*}.\tau)-n.\underline{c}<0$ to $c(n^{*}.\tau)-n.\underline{c}\leq\epsilon\cdot c(n^{*}.\tau)$. In EvalNode, Alg.1 is called with a pruning cost $(1-\epsilon)\cdot c(n^{*}.\tau)$ instead of $c(n^{*}.\tau)$ (Line 2), and the pruning condition is updated to $\mathcal{L}(\pi)\geq(1-\epsilon)\cdot c(n.\tau)$ instead of $\mathcal{L}(\pi)\geq c(n^{*}.\tau)$ (Line 2). Setting $\epsilon=0$ recovers the original, optimal GHOST. Theorem 5 establishes bounded-suboptimality.
@@
GHOST inherently considers all paths, including cyclic ones, as required by rare, complex GCS constraints such as in natarajan2024ixg. Since such cases are uncommon, many GCS shortest-path planners treat cyclic paths as optional and typically restrict to simple paths. Our current implementation similarly restricts path unfolding to simple subpaths between tour endpoints, which limits (bounded-sub)optimality to trajectories in this class (each vertex visited at most $|V|$ times), but yields improved efficiency.
Our implementation sequentially explores tours, fully evaluating all the unfolded paths for one tour before considering the next. While it is possible to compare paths from different tours in parallel, we found it offers little practical benefit, as pruning often eliminates most paths from each tour—especially when $\epsilon$ is used.
Finally, the current LBG aggregates triplet (i.e., 3-tuple) traversal costs, which is the smallest order that provides meaningful lower-bound costs between adjacent sets. Extending to higher-order tuples (e.g., 4-tuples) could yield tighter bounds, but incurs substantial computational overhead with little observed improvement in our experiments.
@@
Greedy (non-optimal) uses greedy search on GCS $G$ with its LBG. Given an initial vertex, a depth-first search chooses the next unvisited vertex $v$ from the current vertex $u$ such that $p=(\cdot,u,v)\in\mathcal{P}$ has the smallest $lb(p)$. After all vertices are visited, an abstract tour is unfolded via Alg. 1, and the search backtracks to explore other tours. The best trajectory found is recorded.
MICP solves GCS-TSP via the unified MICP optimization (marcucci2024graphs), assuming GCS $G=(V,E)$ is complete. For incomplete $G$, $G$ is replaced with $|V|$ copies, denoted as $G^{i},i=1,2,...,|V|$. Each $G^{i}$ represents GCS $G$ at a different step $i$ and is connected to $G^{i+1}$ as described in marcucci2024graphs.
@@
Linear-GCS: We generate incomplete GCS of $M\in[10,50]$ edges. Each convex set $\mathbf{x}_{v}\in\mathcal{X}_{v}\subseteq\mathbb{R}^{2+2}$ comprises of two points $\mathbf{a}_{v},\mathbf{b}_{v}\in\mathbb{R}^{2}$. Each edge $(u,v)$ indicates $\mathcal{X}_{v}\cap\mathcal{X}_{v}\neq\emptyset$ with edge constraint $\mathcal{X}_{e}:=\{(\mathbf{x}_{u},\mathbf{x}_{v})|\mathbf{b}_{u}=\mathbf{a}_{v}\}\subseteq\mathcal{X}_{u}\times\mathcal{X}_{v}$. The trajectory cost $c(\tau)$ is defined as the sum of the Euclidean distance between $\mathbf{a}_{v}$ and $\mathbf{b}_{v}$ for each $v$ along $\tau$.
@@
Ablation Study and Demonstration
@@
Metrics: We compare all methods in terms of solution trajectory cost $c(\tau)$ and runtime. Each method is given a $100$-second limit per random GCS-TSP instance, and metrics for each instance are averaged over $12$ random seeds. We also report the optimality gap $\rho=(c(\tau^{*})-lb^{*})/c(\tau^{*})$ for GHOST and MICP, where $lb^{*}$ is the greatest provable lower bound at termination. For GHOST, $lb^{*}=c(\hat{\pi}_{i})$ is the cost of the next-best abstract tour when $c(\tau^{*})$ is returned. Note that $c(\hat{\pi}_{i})>c(\tau^{*})$ is possible for an optimal $c(\tau^{*})$, so negative $\rho$ values are capped at zero. For MICP, $lb^{*}\leq c(\tau^{*})$ is the current-best relaxed fractional solution cost in the branch-and-bound procedure.
@@
Applications: We demonstrate the versatility of GHOST on Bézier-GCS-TSP instances for coverage, inspection, and task and motion planning. [^1] Check videos at https://sites.google.com/view/ghost-gcs-tsp Fig. 1 and Fig. 4-(c) show continuous trajectories covering all precomputed collision-free convex regions. Fig. 4-(b) shows inspection of a selected subset of the precomputed convex regions, achieved by modifying our RTSP ILP to require only coverage of the subset. Finally, Fig. 4-(d) demonstrates a 7-DoF manipulator (KUKA LBR iiwa) reaching $8$ independent convex sets, each corresponding to a distinct task pose.

### v2

@@ 27       $\text{OPEN}\leftarrow\text{OPEN}\cup\{n_{c}\}$
While the ILP formulation for TSP (miller1960integer) can be adapted to RTSP with a static pairwise distance metric over vertices, this is insufficient for GCS, where edge costs do not provide meaningful estimates or capture the “volume” of tour endpoint sets. Instead, we leverage the triplet-based distance metric established above and reformulate RTSP over all abstract triplets $\hat{p}\in\hat{\mathcal{P}}$. Each $\hat{p}=(u,v,w)$ is formed from every pair of abstract edges $(u,v),(v,w)\in\hat{E}$, with a lower-bound cost defined as $b(\hat{p})=\frac{1}{2}\mathcal{L}(\pi^{*}_{u,v})+b_{\text{mid}}(\hat{p})+\frac{1}{2}\mathcal{L}(\pi^{*}_{v,w})$, where
$b_{\text{mid}}(\hat{p})=\begin{cases}lb_{(u,v,w)},&\text{if }(u,v),(v,w)\in E\\ \min\limits_{p=(u,v,\cdot)\in\mathcal{P}}lb_{p},&\text{if }(u,v)\in E,(v,w)\notin E\\ \min\limits_{p=(v,w,\cdot)\in\mathcal{P}}lb_{p},&\text{if }(u,v)\notin E,(v,w)\in E\\ \min\limits_{p=(\cdot,v,\cdot)\in\mathcal{P}}lb_{p},&\text{if }(u,v),(v,w)\notin E\end{cases}$
[…]
@@
The second inequality follows Theorem 1. Consider $\hat{\pi}=(v_{1},\ldots,v_{k})$ and an unfolded $\pi=\pi_{v_{1},v_{2}}\circ\ldots\circ\pi_{v_{k-1},v_{k}}$, where each subpath $\pi_{v_{i},v_{i+1}}=(v_{i}=w^{(i)}_{1},w^{(i)}_{2},\ldots,w^{(i)}_{\mathcal{K}_{i}}=v_{i+1})$. We decompose $\mathcal{L}(\pi)=\sum_{i=1}^{k}\mathcal{L}(\pi_{v_{i},v_{i+1}})+\sum_{i=1}^{k}lb_{(w^{(i-1)}_{\mathcal{K}_{i-1}-1},v_{i},w^{(i)}_{2})}$, where $v_{k+1}=v_{1}$ (wraparound) and $w^{(0)}_{\mathcal{K}_{0}}=w^{(k-1)}_{\mathcal{K}_{k-1}-1}$ for notation convenience. As each $\mathcal{L}(\pi_{v_{i},v_{i+1}})\geq\mathcal{L}(\pi^{*}_{v_{i},v_{i+1}})$ by definition and $lb_{(w^{(i-1)}_{\mathcal{K}_{i-1}-1},v_{i},w^{(i)}_{2})}\geq b_{\text{mid}}(v_{i-1},v_{i},v_{i+1})$ by construction, summing over all i=1,,ki=1,...,k concludes the proof.
@@
$\sum_{\hat{p}\in\hat{\mathcal{P}}^{\text{ind}}_{e}}y_{\hat{p}}=2,$ $\forall e=(u,v)\in\hat{E}^{+}$ (8)
@@ 25    $\tau\leftarrow$ optimal trajectory conditioned on $\pi$
We introduce a bounded-suboptimal variant, $\epsilon$-GHOST, which trades optimality for computational efficiency. Given a parameter $\epsilon\in[0,1)$, we modify GHOST in two ways: On Line 2, the termination condition is relaxed from $c(n^{*}.\tau)-n.\underline{c}<0$ to $c(n^{*}.\tau)-n.\underline{c}\leq\epsilon\cdot c(n^{*}.\tau)$. In EvalNode, Alg.1 is called with a pruning cost $(1-\epsilon)\cdot c(n^{*}.\tau)$ instead of $c(n^{*}.\tau)$ (Line 2), and the pruning condition is updated to $\mathcal{L}(\pi)\geq(1-\epsilon)\cdot c(n.\tau)$ instead of $\mathcal{L}(\pi)\geq c(n^{*}.\tau)$ (Line 2). Setting $\epsilon=0$ recovers the original, optimal GHOST.
Properties of ($\epsilon$-)GHOST
As in van1999solving; ren2023cbss, the Lawler-Murty procedure at the high level of GHOST guarantees to explore TSP tours in non-decreasing order of their lower-bound costs. Theorem 4 and Corollary 5 establish the (bounded-sub)optimality of ($\epsilon$-)GHOST.
@@
Corollary 5.
GHOST returns an optimal GCS-TSP solution.
Proof 5.
This follows from Theorem 4 with $\epsilon=0.0$.
@@
GHOST inherently considers all paths, including cyclic ones, as required by rare, complex constraints (e.g., natarajan2024ixg). Since such cases are uncommon, many GCS shortest-path planners treat cyclic paths as optional and typically restrict to simple paths. Our implementation similarly restricts path unfolding to simple subpaths between tour endpoints, which limits (bounded-sub)optimality to trajectories in this class (each vertex visited at most $|V|$ times), but yields improved efficiency.
GHOST sequentially explores tours, fully evaluating all the unfolded paths for one tour before considering the next. While it is possible to compare paths from different tours in parallel, we found it offers little practical benefit, as pruning often eliminates most paths from each tour—especially when $\epsilon$ is used.
Finally, extending the current triplet-based LBG and the RTSP ILP to higher-order tuples (e.g., 4-tuples) could yield tighter bounds. However, we observed that this incurs substantial computational overhead with little improvement.
@@
Greedy (non-optimal) uses greedy search on $G$ with its LBG. Given an initial vertex, a depth-first search chooses the next unvisited vertex $v$ from the current vertex $u$ such that $p=(\cdot,u,v)\in\mathcal{P}$ has the smallest $lb_{p}$. After visiting all vertices, an abstract tour is unfolded via Alg. 1, and the search backtracks to explore other tours. The best trajectory found is recorded throughout the greedy search.
MICP solves GCS-TSP via an unified optimization (marcucci2024graphs), assuming $G=(V,E)$ is complete. For incomplete $G$, it is replaced with $|V|$ copies, denoted as $G^{i},i=1,2,...,|V|$. Each $G^{i}$ represents GCS $G$ at a different step $i$ and is connected to $G^{i+1}$ as described in marcucci2024graphs.
@@
Linear-GCS: We generate incomplete GCS of $M\in[10,50]$ edges. Each convex set $\mathbf{x}_{v}\in\mathcal{X}_{v}\subseteq\mathbb{R}^{2+2}$ comprises of two points $\mathbf{a}_{v},\mathbf{b}_{v}\in\mathbb{R}^{2}$. Each edge $(u,v)$ indicates $\mathcal{X}_{u}\cap\mathcal{X}_{v}\neq\emptyset$ with edge constraint $\mathcal{X}_{e}:=\{(\mathbf{x}_{u},\mathbf{x}_{v})|\mathbf{b}_{u}=\mathbf{a}_{v}\}\subseteq\mathcal{X}_{u}\times\mathcal{X}_{v}$. The trajectory cost $c(\tau)$ is defined as the sum of the Euclidean distance between $\mathbf{a}_{v}$ and $\mathbf{b}_{v}$ for each $v$ along $\tau$.
@@
Comparison, Ablation Study, and Demonstration
@@
Metrics: We compare all methods in terms of solution trajectory cost $c(\tau)$ and runtime. Each method is given a $100$-second limit per random GCS-TSP instance, and metrics for each instance are averaged over $12$ random seeds. We also report the optimality gap $\rho=(c(\tau^{*})-lb^{*})/c(\tau^{*})$ for GHOST and MICP, where $lb^{*}$ is the greatest provable lower bound at termination. For GHOST, $lb^{*}=\mathcal{L}(\hat{\pi}_{i})$ is the cost of the next-best abstract tour when $c(\tau^{*})$ is returned. Note that $\mathcal{L}(\hat{\pi}_{i})>c(\tau^{*})$ is possible for an optimal $c(\tau^{*})$, so negative $\rho$ values are capped at zero. For MICP, $lb^{*}\leq c(\tau^{*})$ is the current-best fractional solution cost in branch-and-bound.
@@
Applications: We demonstrate the versatility of GHOST on Bézier-GCS-TSP instances for coverage, inspection, and task and motion planning. Fig. 1 and Fig. 5-(c) show continuous trajectories covering all precomputed collision-free convex regions. Fig. 5-(b) shows inspection of a selected subset of the precomputed convex regions, achieved by modifying our RTSP ILP to require only coverage of the subset. Finally, Fig. 5-(d) demonstrates a 7-DoF manipulator (KUKA LBR iiwa) reaching $8$ independent convex sets, each corresponding to a distinct task pose.